A 50-turn coil has a resistance 2.5 Ω. If it is enclosed in an area of 25 cm2, what would be the rate of change of a magnetic field parallel to its axis to induce a current of 0.5 A in the coil?




1
Expert's answer
2021-12-09T09:56:07-0500

F=BA

where

F=rate of change of magnetic field

B= magnetic density

A= Area

"F=0.0025\times B"

The induced voltage per turn will be the time derivative of that value, so your total induced voltage will is

"V=50×0.0025×\frac{dB}{dt}"

From the formula V=IR we find that V=2.5×0.5=1.25v

Therefore:

"1.25=0.125\frac{dB}{dt}"

"\frac{dB}{dt}=\frac{1.25}{0.125}=10"

So rate of change of magnetic field= 10tesla/sec